OpenStudy (anonymous):

-6x^2=-48 What is "x" using square root terms

OpenStudy (photon336):

1. \[-6x ^{2} = -48\] we've got to isolate x first by dividing both sides by -6 2. \[x ^{2} = \frac{ -48 }{ -6 }\] 3. \[x^{2} = 8\] now we take the square root of both sides because that's the inverse/opposite of x^2 \[\sqrt{x ^{2}} = \sqrt{8}\] this gives us the following \[x = \sqrt{8}\] to simplify the square root of 8, we can try to simplify by doing this. \[\sqrt{8} = \sqrt{4*2} = \sqrt{4}*\sqrt{2} = 2\sqrt{2} = x\]

jhonyy9 (jhonyy9):

@Photon336 --- ATTENTION please ! hope you know that than x^2 = b so than x_1,2 = +/- sqrt(b) - yes ?